OpenStudy (anonymous):

John and Liz are sharing a large pizza. John ate 1/3 of the pizza, and Liz ate 1/5 of the pizza. How much of the pizza is left?

OpenStudy (igreen):

Just add the two fractions. First find a common denominator, which is going to be 15. Convert both fractions to denominators with 15: \(\dfrac{1}{5} \rightarrow \dfrac{3}{15}\) \(\dfrac{1}{3} \rightarrow \dfrac{5}{15}\) Add the numerators and keep the denominator: \(\dfrac{3}{15} + \dfrac{5}{15} \rightarrow \dfrac{3+5}{15} \rightarrow \dfrac{8}{15}\) So there's \(\dfrac{8}{15}\) pizza left.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @nerdychic101